OpenStudy (anonymous):

EVALUATE -3^{2}-(-8)

OpenStudy (anonymous):

Is the question (-3)^2-(-8) or -(3)^2-(-8)?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

It is either 17 or -1 depending on how it is written. (-3)^2-(-8)=17 -(3)^2-(-8)=-1

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

CAN YOU SHOW ME HOW YOU DID THAT